OpenStudy (anonymous):

Find vertex by completing the square. y=-x^2+9x-22

OpenStudy (anonymous):

(4.5,-1.75)

OpenStudy (anonymous):

-x^2+9x=22 ,,, factor out -1 .. -(x^2-9x) =22 ... take half of 9 and square it .. -(x^2-9x+20.25= 22-20.25 ... it's minus 20.25 because of the -1 that was factored out. -(x-9/4)^2 -1.75=0 .. (x=9/4 y=-1.75)

OpenStudy (anonymous):

the easier way is find the axis of symmetry \[x=\frac{ -b }{ 2a }\] plug it in x=\[x=\frac{ -(9) }{ 2(-1) }\] that is how you get the 4.5

OpenStudy (anonymous):

x=4.5 then plug that in y=-x^2+9x-22 y=-4.5^2+9(4.5)-22 that comes up with -1.75

OpenStudy (anonymous):

so that is why the answer is (4.5,-1.75)

OpenStudy (anonymous):

i love questions like this so if you have more just send a message my way